OpenStudy (anonymous):

Identify the local minimum value for the function f(x)= 3x^5 - 5x^3 + 2, using the first derivative test.

OpenStudy (anonymous):

\[f'(x)=15x^4-15x^2\] \[f'(x)=15x^2(x^2-1)\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

critical points at \[\{-1,0,1\}\]

OpenStudy (anonymous):

now \[x^2\geq0\] so the derivative doesn't change sign there. it changes sign at -1 and 1.

OpenStudy (anonymous):

goes from being positive to negative at -1 so function is increasing then decreasing means -1 is a local max. doesn't change sign at 0 so neither max nor min goes from being negative to positive at x = 1 means function is decreasing then increasing so local min

OpenStudy (anonymous):

in sum -1 is a local max 0 is neither max nor min 1 is local min function is increasing on \[(-\infty,-1)\] decreasing on \[(-1,1)\] and then increasing on \[(1,\infty)\]

OpenStudy (anonymous):

clear enough?

OpenStudy (anonymous):

Crystal! :D